OpenStudy (anonymous):

The Minimum Value of : \[(\sin x + cosec x )^2 + (\cos x + \sec x)^2\]

OpenStudy (anonymous):

My work : \[\sin^2 x+ 2 + cosec^2 x + \cos^2x + 2 + \sec^2x\] =\[5 + cosec^2 x + \sec^2 x\] =\[7 + \cot^2x + \tan^2 x\]

OpenStudy (anonymous):

and i am Stuck From here:

OpenStudy (anonymous):

5 + 1/sin^2 x + 1/cos^2 x = 5 + 1/sin^2 x cos^2 x = 5+ 2/sin (2x)

OpenStudy (anonymous):

That...was a gud..Idea..) then

OpenStudy (anonymous):

now the minimum occurs where sin 2x is at maximum isnt it?

OpenStudy (anonymous):

Yes...)

OpenStudy (anonymous):

so......7

OpenStudy (anonymous):

thxxx...).....that was awesome..brother

OpenStudy (anonymous):

welcme

OpenStudy (anonymous):

WELL THE MINIMUM of 5+ 2/sin (2x) is 3 when sin 2x =-1

OpenStudy (anonymous):

Isnt it @Yahoo! and @him1618

OpenStudy (anonymous):

For the eq to be Min the Denominator Shuld be Maximum

OpenStudy (anonymous):

MAx Value of Sin =1